OpenStudy (anonymous):

What is the method for factoring -x^3+3x+2?

OpenStudy (somethingawesome):

One way is to find a root, then factor out that term. We can make an educated guess using the Rational Root Theorem, but however we do it, notice that -2^3+3(2)+2 = -8+6+2 = 0, so x=2 is a root. That means the term (x-2) factors out. Do you know how to divide polynomials? Because now we want to divide -x^3+3x+2 by x-2 to get what the other terms are. What's left is a quadratic, so we can factor that just fine.

OpenStudy (anonymous):

Is there a systematic way of solving it?

OpenStudy (dumbcow):

I would say that is the systematic way..try to factor out (x +- 1) or (x+- 2) using polynomial division or synthetic division.

OpenStudy (anonymous):

I don't think I understand. Can you show me with this equation?

OpenStudy (dumbcow):

somethingawesome already showed you that (x-2) does indeed factor out evenly. But the way to approach it is there are only 4 possible terms which will factor out evenly... (x +- 1) or (x +- 2). do you know how to do synthetic division?

OpenStudy (somethingawesome):

If you want an equation like the quadratic one but for cubics, there are some here: http://en.wikipedia.org/wiki/Cubic_function I don't think they'll be very helpful though.

OpenStudy (dumbcow):

2| -1 0 3 2 -2 -4 -2 _______________ -1 -2 -1 0 This is an example of dividing by (x-2) using synthetic division. The last number is zero which is the remainder, thus it divides evenly. The other numbers are the coefficients in the quotient Factored form: (x-2)(-x^2 -2x -1) Pull out a neg -(x-2)(x^2 +2x +1) factor quadratic...notice its a perfect square -(x-2)(x+1)^2

OpenStudy (anonymous):

Got it. Thanks!

OpenStudy (dumbcow):

ok do you understand the synthetic division?

OpenStudy (anonymous):

I can do regular polynomial divsion, and I'm looking up how to do synthetic division on purplemath. I guess if worst comes to worst I can always use long division

OpenStudy (dumbcow):

no either way is fine..synthetic is just easier at times especially in these cases the goal is just to find a root and divide it out to get it into factored form. you don't need to spend too much time learning it if you are comfortable with regular division